OpenStudy (goformit100):

Resolve into factors : 99899

OpenStudy (goformit100):

@bahupepe @mayankdevnani @hartnn @UnkleRhaukus

OpenStudy (anonymous):

\[99899 =283 · 353\]

sam (.sam.):

Prime factors?

OpenStudy (goformit100):

@madmaddiemac @math0101 @math111

sam (.sam.):

1\283\ 353\ 99899

OpenStudy (anonymous):

got it

OpenStudy (unklerhaukus):

what technique have you used @bahupepe ?

hartnn (hartnn):

even wolf can tell us that. explain how you got it ?

OpenStudy (goformit100):

I solved like this

OpenStudy (goformit100):

99899 = f (10) Then f(10) = g (10) . h(10) if So, = = g (x) . h (x) Hence 99899 = f(10) = g(10) x h(10) = 283 x 353

OpenStudy (goformit100):

is it correct ?

hartnn (hartnn):

lol

hartnn (hartnn):

where are the steps in between ?

OpenStudy (unklerhaukus):

haha

sam (.sam.):